OpenStudy (anonymous):

x^(2)+5x-7=0 the exact solutions are x=____?_____ the approximate solutions to three decimal places are x=____?____

OpenStudy (vishal_kothari):

-6.140054945 and 1.1400549

OpenStudy (anonymous):

Use the quadratic formula.

OpenStudy (bahrom7893):

hmm vishal those are neither exact, nor approximate to 3 decimals lol

OpenStudy (anonymous):

http://edu.glogster.com/media/5/21/8/90/21089020.jpg

OpenStudy (bahrom7893):

D = 25 + 28 = 53 x1 = (-5-sqrt(53))/2 x2 = (-5+sqrt(53))/2

OpenStudy (bahrom7893):

Those are exact

OpenStudy (vishal_kothari):

they are exact...

OpenStudy (bahrom7893):

not really lol

OpenStudy (anonymous):

this is what i came up with am I right???? the exact solutions are x=-5+sqr53/2,-5-5sqr53/2 the approximate solutions to three decimal places are x= 1.1401,-6.1401

OpenStudy (anonymous):

@bahrom's answer IS exact.

OpenStudy (bahrom7893):

crzy those are 4 decimals 3 would be: 1.140 and -6.140

OpenStudy (vishal_kothari):

yeah..

OpenStudy (anonymous):

@vishal kothari 's is approx.

OpenStudy (bahrom7893):

or technically 1.14 or -6.14

OpenStudy (vishal_kothari):

kk